OpenStudy (anonymous):

what is the domain of y=x*sq rt of (2-x^2) and why?

OpenStudy (science0229):

First of all, whatever is inside the square root has to be greater of equal to 0, right?

OpenStudy (science0229):

So, 2-x^2>=0 x^2<=2 -sqrt(2)<=x<=sqrt(2)

OpenStudy (anonymous):

i do not understand how you got that last line

OpenStudy (anonymous):

i got x<= positive and negative sqrt of 2

OpenStudy (science0229):

Do you know how to graph y=x^2?

OpenStudy (science0229):

and y=2?

OpenStudy (anonymous):

yes

OpenStudy (science0229):

Can you draw both of the graphs on the same coordinate plane for me?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):